# Matchings in Matroids over Abelian Groups, III

### v1

@@ 3.2. Matching paving matroids over abelian groups (asymmetric case)
We now extend this result to general paving matroids by incorporating the hyperplane-nullity parameter. In particular, we generalize the sparse paving case ($t\leq 1$) to arbitrary paving matroids, with a sharp dependence on the parameter $t$.
@@
Let $\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ be a basis of $M$, and define $b_{i}:=f(a_{i})$ for allb $i\in[n]$. Set $\mathcal{N}=\{b_{1},\ldots,b_{n}\}\subseteq E(N)$. Since $N$ is a paving matroid of rank $n$, the set $\mathcal{N}$ is either a basis or a circuit.
@@
There exist distinct elements $x_{1},\dots,x_{t}\in E(N)\setminus\mathcal{N}$ and distinct indices $i_{1},\dots,i_{t}\in[n]$ such that
$a_{i_{j}}+x_{j}\notin E(M)\quad\text{for every }j\in[t].$
@@
We proceed by induction on $t$.
Base case ($t=1$). We show that there exist $x\in E(N)\setminus\mathcal{N}$ and $a\in\mathcal{M}$ such that $a+x\notin E(M)$. Suppose, for contradiction, that no such pair $(a,x)$ exists. Then
@@
Inductive hypothesis. Assume that distinct elements $x_{1},\dots,x_{t-1}\in E(N)\setminus\mathcal{N}$ and distinct indices $i_{1},\dots,i_{t-1}\in[n]$ have been chosen such that
$a_{i_{j}}+x_{j}\notin E(M)\quad\text{for every }j\in[t-1].$
Inductive step. We show that there exist
$x_{t}\in E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{t-1}\})\quad\text{and}\quad i_{t}\in[n]\setminus\{i_{1},\dots,i_{t-1}\}$
@@
(5) $(E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{t-1}\}))+(\mathcal{M}\setminus\{a_{i_{1}},\dots,a_{i_{t-1}}\})\subseteq E(M).$
Applying Theorem 2.5 to the sets $\mathcal{M}\setminus\{a_{i_{1}},\dots,a_{i_{t-1}}\}$ and $E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{t-1}\})$, we obtain a subgroup $H\leq G$ such that
(6) $\Big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\}))\Big|$
$\qquad\geq|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})|-|H|,$
$(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\}))+H$
$\qquad=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})).$
@@
(7) $|E(M)|$ $\geq\big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\}))\big|$
$\geq|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})|-1$
$=|E(N)|-2t+1.$
@@
$\Big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\})+\big(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})\big)\Big|\\ =\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})\big|-1.$
@@
$\big(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\},\,E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})\big)$
@@
  • $|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}|=n-(t-1)>1$;
@@
    $|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})|=|E(N)|-(n+t-1)=|E(M)|+2t-1-(n+t-1)>t>1;$
  • $|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})|-1=|E(M)|\leq|E(N)|-1\leq p(G)-2.$
@@
$E(M)=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\})+\big(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})\big)$
@@
$|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})|-1=|E(M)|,$
@@
$|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})|-1=|E(M)|-1.$
@@
    $\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})\big|-1=|E(M)|,$
@@
    $\big(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\},\,E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})\big)$
    is critical and satisfies the conditions of Theorem 2.6. Hence both $\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}$ and $E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})$ are progressions with the same common difference. Since
    $E(M)=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})),$
@@
    $\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{t-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{t-1}\})\big|-1=|E(M)|-1,$
@@
Note that Theorem 3.8 does not address the case $t=0$, since this forces NN to be a uniform matroid (see Corollary 2.4). Matchability in uniform matroids is straightforward and has been completely characterized in [4, Proposition 2.9].
@@ 3.3. Matchable bases through stressed hyperplanes
When matchings between matroid bases are not available, stressed hyperplanes show that compatibility can often be recovered by relaxing independence constraints. A stressed hyperplane may be viewed as a minimal obstruction to uniformity. Relaxing such hyperplanes, turning certain dependent sets into bases, gradually transforms a paving matroid toward a uniform matroid while preserving structure. Theorem 3.20 shows that this process offers an alternative route to matchability: even if bases of $M$ cannot be matched to those of $N$ directly, they may be matched to bases of a relaxation $\mathrm{Rel}_{S}(N)$. Thus matchability is not a fragile property tied to exact independence, but a stable phenomenon that persists under natural modifications. Relaxation therefore, links the restrictive setting of paving matroids with the uniform case, where matching problems are straightforward.
@@ 3.4. Concluding remarks and future directions
    Assuming the conditions of Theorem 3.20, what is the minimal number of steps in this hierarchy needed to guarantee that $M$ can be matched to some consecutive relaxations of $N$? Clearly this process will be terminated as in the worst case scenario $M$ will be eventually matched to $U_{n,m}$ Is a single relaxation always sufficient, or can certain pairs $(M,N)$ require multiple relaxations within the hierarchy?
Data sharing: Data sharing is not applicable to this article as no datasets were generated or analyzed.
Conflict of interest: To our best knowledge, no conflict of interests, whether of financial or personal nature, has influenced the work presented in this article.

### v2

@@ 3.2. Matching Paving Matroids over Abelian Groups (Asymmetric Case)
We now extend this result to general paving matroids by incorporating the hyperplane-nullity parameter. In particular, we generalize the sparse paving case ($t\leq 1$) to arbitrary paving matroids, with explicit dependence on the parameter $t$.
@@
Let $\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ be a basis of $M$, and define $b_{i}:=f(a_{i})$ for all $i\in[n]$. Set $\mathcal{N}=\{b_{1},\ldots,b_{n}\}\subseteq E(N)$. Since $N$ is a paving matroid of rank $n$, the set $\mathcal{N}$ is either a basis or a circuit.
@@
For every integer $s$ with $s\in[t]$, there exist distinct elements $x_{1},\dots,x_{s}\in E(N)\setminus\mathcal{N}$ and distinct indices $i_{1},\dots,i_{s}\in[n]$ such that
$a_{i_{j}}+x_{j}\notin E(M)\qquad\text{for every }j\in[s].$
In particular, the conclusion holds for $s=t$.
@@
We proceed by induction on $s$.
Base case ($s=1$). We show that there exist $x\in E(N)\setminus\mathcal{N}$ and $a\in\mathcal{M}$ such that $a+x\notin E(M)$. Suppose, for contradiction, that no such pair $(a,x)$ exists. Then
@@
Inductive hypothesis. Assume that distinct elements $x_{1},\dots,x_{s-1}\in E(N)\setminus\mathcal{N}$ and distinct indices $i_{1},\dots,i_{s-1}\in[n]$ have been chosen such that
$a_{i_{j}}+x_{j}\notin E(M)\quad\text{for every }j\in[s-1].$
Inductive step. We show that there exists
$x_{s}\in E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{s-1}\})\quad\text{and}\quad i_{s}\in[n]\setminus\{i_{1},\dots,i_{s-1}\}$
@@
(5) $(E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{s-1}\}))+(\mathcal{M}\setminus\{a_{i_{1}},\dots,a_{i_{s-1}}\})\subseteq E(M).$
Applying Theorem 2.5 to the sets $\mathcal{M}\setminus\{a_{i_{1}},\dots,a_{i_{s-1}}\}$ and $E(N)\setminus(\mathcal{N}\cup\{x_{1},\dots,x_{s-1}\})$, we obtain a subgroup $H\leq G$ such that
(6) $\Big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\}))\Big|$
$\qquad\geq|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-|H|,$
$(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\}))+H$
$\qquad=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})).$
@@
(7) $|E(M)|$ $\geq\big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\}))\big|$
$\geq|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1$
$=|E(N)|-2s+1\geq|E(N)|-2t+1.$
@@
$\Big|(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+\big(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)\Big|\\ =\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big|-1.$
@@
$\big(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\},\,E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)$
@@
  • $|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|=n-(s-1)\geq n-(t-1)>1$;
@@
    $|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|$ $=|E(N)|-(n+s-1)$
    $\geq|E(N)|-(n+t-1)$
    $=|E(M)|+2t-1-(n+t-1)$
    $>t>1.$
  • $|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|$ $+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1$
    $=|E(M)|$
    $\leq|E(N)|-1$
    $\leq p(G)-2.$
@@
$E(M)=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+\big(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)$
@@
$|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1=|E(M)|,$
@@
$|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}|+|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})|-1=|E(M)|-1.$
@@
    $\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big|-1=|E(M)|,$
@@
    $\big(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\},\,E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big)$
    is critical and satisfies the conditions of Theorem 2.6. Hence both $\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}$ and $E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})$ are progressions with the same common difference. Since
    $E(M)=(\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\})+(E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})),$
@@
    $\big|\mathcal{M}\setminus\{a_{i_{1}},\ldots,a_{i_{s-1}}\}\big|+\big|E(N)\setminus(\mathcal{N}\cup\{x_{1},\ldots,x_{s-1}\})\big|-1=|E(M)|-1,$
@@
Note that Theorem 3.10 does not address the case $t=0$, since this forces NN to be a uniform matroid (see Corollary 2.4). Matchability in uniform matroids is straightforward and has been completely characterized in [6, Proposition 2.9].
Remark 3.18.
Note that in the asymmetric problem of the sparse paving case, dependence is controlled by circuit-hyperplanes, whereas in the general paving case, the size of the obstruction is measured by hyperplane nullity. Theorem 3.10 shows that the existence of matchings is governed not merely by the size gap |E(N)||E(M)||E(N)|-|E(M)|, but also by the extent to which $N$ deviates from uniformity, as quantified by the hyperplane nullity parameter $t$.
@@ 3.3. Matchable Bases through Stressed Hyperplanes
Let $E$ be a nonempty set. Recall that A matroid basis system is a nonempty family $\mathcal{B}\subseteq 2^{E}$ of subsets, called bases, that satisfies the exchange axiom: for all distinct $B,B^{\prime}\in\mathcal{B}$ and every $x\in B\setminus B^{\prime}$, there exists $y\in B^{\prime}\setminus B$ such that $(B\setminus\{x\})\cup\{y\}\in\mathcal{B}$. It follows that $\mathcal{B}$ is the set of bases of a matroid.
@@
When matchings between matroid bases are not available, stressed hyperplanes show that compatibility can often be recovered by relaxing independence constraints. A stressed hyperplane may be viewed as a minimal obstruction to uniformity. Relaxing such hyperplanes, turning certain dependent sets into bases, gradually transforms a paving matroid toward a uniform matroid while preserving structure. Theorem 3.23 shows that this process offers an alternative route to matchability: even if bases of $M$ cannot be matched to those of $N$ directly, they may be matched to bases of a relaxation $\mathrm{Rel}_{S}(N)$. Thus matchability is not a fragile property tied to exact independence, but a stable phenomenon that persists under natural modifications. Relaxation therefore links the restrictive setting of paving matroids with the uniform case, where matching problems are straightforward.
@@ 3.4. Concluding Remarks and Future Directions
    Assuming the conditions of Theorem 3.23, what is the minimal number of steps in this hierarchy needed to guarantee that $M$ can be matched to some consecutive relaxations of $N$? Clearly, this process must terminate, since in the worst-case scenario $M$ is eventually matched to $U_{n,m}$. Is a single relaxation always sufficient, or can certain pairs $(M,N)$ require multiple relaxations within the hierarchy?
Data sharing. Data sharing is not applicable to this article as no datasets were generated or analyzed.
Conflict of interest. To the best of our knowledge, no conflict of interest, whether of a financial or personal nature, has influenced the work presented in this article.